The perimeter of a rectangular backyard is 6x + 6 yards. If the width is x yards, find a binomial that represents the length
Answer by mananth(16946)

2(L+W)=perimeter = 6x+6
L+W = 6x+6 /2
L+x= (6x+6)/2
L=(6x+6)/2 -x
LCD =2
L=((6x+6)-2x) /2
L=(6x+6-2x) /2
L= (4x+6) /2
L=2x+3
...